OpenStudy (anonymous):

If \(\Large x^2=8x+y\) and \(\Large y^2=x+8y\), determine the numerical value of \(\Large x^2+y^2\). Assume \(\Large x\neq y\).

ganeshie8 (ganeshie8):

\[\large x^2 + y^2 = 9(x+y)\]

OpenStudy (zzr0ck3r):

says numerical

OpenStudy (anonymous):

so, what is the numerical value of it ?

ganeshie8 (ganeshie8):

\[\large x^2 - y^2 = 7(x-y) \implies x+y = 7 ~~~\because x\ne y \]

ganeshie8 (ganeshie8):

plug that in first equation ^^

OpenStudy (anonymous):

Or multiply x^2+y^2 by x^2-y^2

ganeshie8 (ganeshie8):

Oh yes thats even better :)

OpenStudy (anonymous):

:)

OpenStudy (zzr0ck3r):

this is not my best area can you explain this step @ganeshie8 \(\large x^2 - y^2 = 7(x-y) \implies x+y = 7 ~~~\because x\ne y\)

OpenStudy (zzr0ck3r):

nm difference is squares.... sorry

ganeshie8 (ganeshie8):

yes :) just cancelled out the nonzero factor (x-y)

OpenStudy (zzr0ck3r):

very clever